There are six people, six seats. What is the probability that both the youngest girl and the eldest boy (3 girls, 3 boys in total) are seated at each end?
(yes, this one is easy...but the next question i don't get at all...)
.
.
.
What is the probability that exactly one of them (not at least, but exactly one of them) gets seated at the end?
Thanks for any help at all!
: There are six people, six seats. What is the probability that both the youngest girl and the eldest boy (3 girls, 3 boys in total) are seated at each end?
(yes, this one is easy...but the next question i don't get at all...)
.
.
.
What is the probability that exactly one of them (not at least, but exactly one of them) gets seated at the end?
Thanks for any help at all!

Answer by stanbon(19743)

There are six people, six seats. What is the probability that both the youngest girl and the eldest boy (3 girls, 3 boys in total) are seated at each end?
(yes, this one is easy...but the next question i don't get at all...)
# of ways to place the youngest and eldest as described: 2
# of ways to arrange the remaining 4: 4! 24
---------
Total # of ways to get the pattern you want: 2*24 = 48
Total # of ways to arrange the 6 with no restrictions: 6! = 720
----------
Probability of succeeding : 48/720 = 0.066667
========================================================.
.
.
What is the probability that exactly one of them (not at least, but exactly one of them) gets seated at the end?
# of ways to get one of them on the end: 2*2 = 4
# of ways to place the one not on the end: 4
# of ways to place the remaining 4: 4! = 24
Total # of ways to succeed: 4*4*24 = 96
Total # of ways to arrange the 6 with no restrictions: 72
Probability of succeeding: 96/720 = 0.133333..
Cheers,
Stan H.
------------
# of ways